If S={x | 0<x<12}, M={x | 2<x<6},and N={x | 4<x<12}, find


(a) M∪N (b) M∩N (c) (S∩N)′ (d) M′∩N′

a) x | 2 < x <12

b) x | x = 5

c) little confused on

d)also confused on, Help please.
Answer #1

C) (S \cap N) = {x | 4 < x < 12}

(S \cap N)' = {x | -\infty < x < 4 or, 12 < x < \infty}

D) (M U N) = {x | 2 < x < 12}

(M U N)' = (M' \cap N') = {x | -\infty < x < 2 or, 12 < x < \infty}